# An Auditable AI Agent Loop for Empirical Economics: A Case Study in Forecast Combination

### 18 Mar 2026 v1

@@ Abstract
AI coding agents make empirical specification search fast and cheap, but they also widen hidden researcher degrees of freedom. Building on an open-source agent-loop architecture, this paper recasts a minimal coding loop as a transparent protocol for empirical economics. In a forecast-combination illustration, multiple independent agent runs outperform standard benchmarks in the original rolling evaluation, but not all continue to do so on a post-search holdout. Logged search and holdout evaluation together make adaptive specification search visible and help distinguish robust improvements from sample-specific discoveries.
Keywords: Agent loops, Specification search, Researcher degrees of freedom, Responsible AI, Forecast combination
@@ 1 Introduction
This paper studies one response. Building on Karpathy2026’s (Karpathy2026) open-source autoresearch, I recast a simple agentic coding loop as a research protocol for empirical economics. The protocol has four ingredients: a written instruction contract, an immutable evaluator, a single editable script, and a complete experiment log. Together they do not eliminate adaptive search; they make it inspectable, bounded, and easier to disclose. The contribution is not a new agent architecture, but a way to use such loops in empirical work without treating the search path as invisible.
@@
That split result is the main message. The experiment logs show that nominal hyperparameter tuning readily drifts into broader method discovery. Some departures survive external evaluation; others do not. The value of the protocol is therefore twofold: it records how the agent searched, and it allows the resulting departures to be judged rather than hidden. For empirical economics, that is the relevant promise of agentic search: not automation without discretion, but automation with an audit trail.
Recent work on AI for science spans a wide range of autonomy. On the code and data-analysis side, systems such as AlphaEvolve (NovikovEtAl2025), DS-STAR (ChenEtAl2025), and related work on expert-level empirical software (AygunEtAl2025) use large language models to iteratively rewrite code, plan analyses, or refine scientific software against automated evaluators or verification steps. At the other extreme, systems such as The AI Scientist (LuEtAl2024) and AI co-scientist (GottweisEtAl2025) move toward broader research automation, spanning idea generation, experiment design, analysis, and scientific communication. In economics, Korinek2025 surveys the use of AI agents across research tasks, while DawidEtAl2025 develop broader agentic workflows for economic research. Relative to this literature, our contribution is deliberately narrower: we study evaluator-locked local search inside an existing empirical workflow, where the researcher fixes the question, data, evaluator, editable surface, and search budget, and the agent is autonomous only within those boundaries. The goal is not full research automation, but an auditable protocol that can be integrated into standard empirical practice.
@@ 2 An auditable agent-loop protocol
Consider a researcher who starts with a baseline empirical model and wants to explore nearby alternatives using an AI coding agent. [^1] By AI coding agent we mean a general-purpose coding assistant such as Claude Code or OpenAI Codex that plans, writes code, and executes tool calls within a development environment. We formalize the search as follows. Let $C$ denote a written instruction contract that specifies the research objective, admissible modifications, and search budget. Let $\mathcal{T}(C)$ denote the class of admissible editable scripts under $C$, and let $S(\tau;D)$ denote the scalar score returned by an immutable evaluator when script $\tau\in\mathcal{T}(C)$ is run on data $D$. Both $\mathcal{T}(C)$ and $S$ are fixed ex ante by the researcher as part of the research design. In practice, the contract fixes the evaluator, editable surface, and no-look-ahead constraint, but it need not impose a machine-checkable restriction that candidate scripts belong to a particular model family; the effective admissible set $\mathcal{T}(C)$ may therefore be broader than the researcher’s semantic intent. A complete experiment log $L$ records every attempted modification and its outcome. These four elements—contract, evaluator, editable script, and log—constitute the protocol.
When lower scores are preferred, the research problem is
$\tau^{\star}\in\arg\min_{\tau\in\mathcal{T}(C)}S(\tau;D).$
@@
$\hat{\tau}_{K}\in\arg\min_{\tau\in\{\tau_{0},\tau_{1},\ldots,\tau_{K}\}}S(\tau;D).$
This distinction matters. $\tau^{\star}$ is the best admissible script in the full class $\mathcal{T}(C)$. $\hat{\tau}_{K}$ is the best script found by a particular bounded search. The gap between the two reflects the cost of finite search. The audit log records the outcome of every experiment as an ordered sequence $L_{K}=\bigl((k,\tau_{k},s_{k},d_{k})\bigr)_{k=0}^{K}$, where $s_{k}=S(\tau_{k};D)$ when the run succeeds, $s_{k}=+\infty$ when it crashes, and $d_{k}\in\{\textit{keep},\,\textit{discard},\,\textit{crash}\}$ records the outcome status. Algorithm 1 summarizes the protocol.
@@
1:Instruction contract $C$, evaluator $S(\cdot\,;D)$, initial script $\tau_{0}$, budget $K$
@@
9:Best script $\hat{\tau}_{K}=\tau^{\mathrm{best}}$ and complete log $L_{K}$
The protocol does not impose a model of how the sequence of candidates is generated; that depends on the underlying large language model and the coding environment. Because the agent observes what it has already tried, the search is path dependent and stochastic: independent runs from the same starting point will generally follow different paths and may reach different solutions. The search is adaptive but bounded. The agent searches freely within the editable script, but the evaluator, data, and scoring rule are fixed ex ante. This separation between immutable research design and editable implementation is analogous to a pre-analysis plan, except that the audit trail is generated automatically and covers every specification the agent considered. Implementation details, including the instruction contract, are in the Appendix.
@@ 3.2 Search-sample and holdout results
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Best avg flex and best $(N,W)$ avg are from DieboldShin2019, Tables 4 and 5. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2025); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) uses $\lambda$ selected on the search sample for all windows; peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately. Full results including intermediate methods are in Appendix E.
@@ 3.3 What the log reveals
The instruction contract asks the agent to tune regularization parameters for peLASSO. The experiment logs [^2] Appendix C reproduces an excerpt from the experiment log of Run 3, and Appendix D gives a detailed description of the method discovered by each run., however, reveal that two of the three runs drifted beyond that task: Run 1 toward stability selection and performance weighting, and Run 2 toward a ranking based method with bias correction. Run 3 stayed closest to the original task, using adaptive LASSO with forward cross validation for lambda selection, though it added an egalitarian elastic net blend. The holdout makes these departures evaluable. Run 2’s bias correction, which exploits serial correlation in combined forecast errors, turns out to generalize and accounts for most of its holdout gain. Run 3’s egalitarian elastic net blend, which drove much of its search-sample improvement, degrades holdout performance. The protocol does not prevent the agent from leaving the intended design space; it ensures that every departure is recorded, so that a holdout or other external evidence can later distinguish structural insight from sample adaptation.
@@ 4 Conclusion
Agent loops can widen researcher degrees of freedom, but they need not do so invisibly. This paper recasts a minimal agent loop architecture as an auditable protocol for agentic specification search, requiring only a fixed evaluator, a single editable script, a finite search budget, and a complete experiment log. In the forecast combination illustration based on DieboldShin2019, three independent runs improve on the search sample, but a genuine holdout shows that only two of the three discovered methods generalize. The contribution of the protocol is not to eliminate adaptive search, but to make it visible, bounded, and evaluable. Because the protocol disciplines the search process rather than the subject matter, it applies wherever automated specification search is used. Appendix F sketches how the same architecture maps to regression specification search, structural VAR identification, treatment effect estimation, and other standard empirical tasks.
Four limitations deserve emphasis. First, a language model may encode background knowledge that overlaps with the holdout period, so residual look-ahead concerns remain. Second, the model’s internal trade-offs between score improvement and contract boundaries are opaque; it is often unclear how those objectives are balanced. Third, the effective search class may recombine existing ideas more readily than it generates novel ones. Fourth, the current search over $\mathcal{T}(C)$ is greedy: propose one candidate, evaluate, keep or discard, repeat. More structured strategies such as the evolutionary population-based search in AlphaEvolve (NovikovEtAl2025) could explore the space more efficiently.
@@ A.1 Four-file architecture
program.md is a plain-text instruction contract that defines the research objective, the admissible modifications, the rules, and the experiment loop of Algorithm 1. train.R is the editable script containing one function the agent rewrites. prepare.R is the immutable evaluator: it loads the data, sources train.R, and returns the scalar score. results.tsv is a tab-separated audit log in which each row records a commit identifier, the score, the outcome status, and a verbal description of the strategy attempted. In the working directory, the git commit history preserves the actual code of every retained $\tau_{k}$. [^3] The replication archive includes the TSV logs and final code but not the full git object store. Full recoverability from commit history requires access to the original working directory.
@@ E.3 Full holdout results
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2025); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) and best individual use parameters selected on the search sample and are evaluated without re-optimization on the holdout. peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately.
@@ E.5 Independent dataset robustness
  • Notes: Same as Table 4 but using the independently constructed dataset (all rows from downloaded SPF and Eurostat data, no verbatim copy from the original H1_gdp.csv). Search-sample RMSEs differ from Table 4 due to GDP vintage differences and independent imputation. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test using the EWC fixed-$b$ approximation (ShinSchor2025); see Table 4 notes for details and caveats. peLASSO ex post (original) and best individual use parameters selected on the search sample; see Table 4 notes for details.

### 19 Mar 2026 v2

@@ Abstract
AI coding agents make empirical specification search fast and cheap, but they also widen hidden researcher degrees of freedom. Building on an open-source agent-loop architecture, this paper adapts that framework to an empirical economics workflow and adds a post-search holdout evaluation. In a forecast-combination illustration, multiple independent agent runs outperform standard benchmarks in the original rolling evaluation, but not all continue to do so on a post-search holdout. Logged search and holdout evaluation together make adaptive specification search more transparent and help distinguish robust improvements from sample-specific discoveries.
Keywords: Agent loops, Autoresearch, Specification search, Researcher degrees of freedom, Responsible AI, Forecast combination
@@ 1 Introduction
This paper studies one response. Building on Karpathy2026’s (Karpathy2026) open-source autoresearch, I adapt a simple agentic coding loop to a research protocol for empirical economics. The protocol has four ingredients: a written instruction contract, an immutable evaluator, a single editable script, and a complete experiment log. I then add an explicit post-search holdout evaluation. Together, they do not eliminate adaptive search, but they make it inspectable, bounded, and easier to disclose, while also making validation beyond the search sample explicit. The point is therefore not a new agent architecture, but a way to use an existing one in empirical work with clearer external validation.
@@
That split result is the main message. The experiment logs show that nominal hyperparameter tuning readily drifts into broader method discovery. Some departures survive external evaluation; others do not. The value of the framework is therefore twofold: the underlying loop records how the agent searched, and the added holdout evaluation allows the resulting departures to be judged rather than hidden. For empirical economics, that is the relevant promise of agentic search: not automation without discretion, but automation with an audit trail and explicit post-search holdout evaluation.
Recent work on AI for science spans a wide range of autonomy. On one end, systems such as AlphaEvolve (NovikovEtAl2025) and DS-STAR (ChenEtAl2025) use large language models to iteratively rewrite code and refine analyses against automated evaluators. On the other end, systems such as The AI Scientist (LuEtAl2024) and AI co-scientist (GottweisEtAl2025) pursue broader research automation across idea generation, experiment design, analysis, and communication. In economics, Korinek2025 surveys AI agents for research, while DawidEtAl2025 develop broader agentic workflows. Our contribution is narrower: we study evaluator-locked local search within a fixed empirical workflow, where the researcher sets the question, data, evaluator, editable surface, and search budget. The goal is not full automation, but an auditable protocol for standard empirical practice.
@@ 2 An auditable agent-loop protocol
Consider a researcher who starts with a baseline empirical model and wants to explore nearby alternatives using an AI coding agent. [^1] By AI coding agent we mean a general-purpose coding assistant such as Claude Code or OpenAI Codex that plans, writes code, and executes tool calls within a development environment. We formalize the search as follows. Let $C$ denote a written instruction contract that specifies the research objective, admissible modifications, and search budget. Let $\mathcal{T}(C)$ denote the class of admissible editable scripts under $C$. Let the available data be partitioned as $D=(D^{S},D^{H})$, where $D^{S}$ is the search sample and $D^{H}$ is a post-search holdout reserved for external validation. Let $S(\tau;D^{S})$ denote the scalar score returned by an immutable evaluator when script $\tau\in\mathcal{T}(C)$ is run on search sample $D^{S}$. Both $\mathcal{T}(C)$ and $S$ are fixed ex ante by the researcher as part of the search design. In practice, the contract fixes the evaluator, editable surface, and no-look-ahead constraint, but it need not impose a machine-checkable restriction that candidate scripts belong to a particular model family; the effective admissible set $\mathcal{T}(C)$ may therefore be broader than the researcher’s semantic intent. A complete experiment log $L$ records every attempted modification and its outcome. These four elements—contract, evaluator, editable script, and log—constitute the protocol.
When lower scores are preferred, the search problem is
$\tau^{\star}\in\arg\min_{\tau\in\mathcal{T}(C)}S(\tau;D^{S}).$
@@
$\hat{\tau}_{K}\in\arg\min_{\tau\in\{\tau_{0},\tau_{1},\ldots,\tau_{K}\}}S(\tau;D^{S}).$
This distinction matters. $\tau^{\star}$ is the best admissible script in the full class $\mathcal{T}(C)$. $\hat{\tau}_{K}$ is the best script found by a particular bounded search. The gap between the two reflects the cost of finite search. The audit log records the outcome of every experiment as an ordered sequence $L_{K}=\bigl((k,\tau_{k},s_{k},d_{k})\bigr)_{k=0}^{K}$, where $s_{k}=S(\tau_{k};D^{S})$ when the run succeeds, $s_{k}=+\infty$ when it crashes, and $d_{k}\in\{\textit{keep},\,\textit{discard},\,\textit{crash}\}$ records the outcome status. Algorithm 1 summarizes the protocol.
@@
1:Instruction contract $C$, evaluator $S(\cdot\,;D^{S})$, search sample $D^{S}$, initial script $\tau_{0}$, budget $K$
@@
The protocol does not impose a model of how the sequence of candidates is generated; that depends on the underlying large language model and the coding environment. Because the agent observes what it has already tried, the search is path dependent and stochastic: independent runs from the same starting point will generally follow different paths and may reach different solutions. The search is adaptive but bounded. The agent searches freely within the editable script, but the evaluator, search sample, and scoring rule are fixed ex ante. This separation between immutable search design and editable implementation is analogous to a pre-analysis plan, except that the audit trail is generated automatically and covers every specification the agent considered. Implementation details, including the instruction contract, are in the Appendix.
Once the search is complete, the researcher may evaluate the selected script $\hat{\tau}_{K}$ on the holdout $D^{H}$. Let $S^{H}(\tau;D^{H})$ denote a scoring function computed on the holdout. The realized holdout score $S^{H}(\hat{\tau}_{K};D^{H})$ provides an external check on whether improvements discovered during adaptive search on $D^{S}$ generalize beyond the search sample. Because $D^{H}$ need not even exist during the search, the holdout evaluation is a researcher action, not part of the agentic loop.
@@ 3.2 Search-sample and holdout results
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Best avg flex and best $(N,W)$ avg are from DieboldShin2019, Tables 4 and 5. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) uses $\lambda$ selected on the search sample for all windows; peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately. Full results including intermediate methods are in Appendix E.
@@ 3.3 What the log reveals
The instruction contract asks the agent to tune regularization parameters for peLASSO. The experiment logs [^2] Appendix C reproduces an excerpt from the experiment log of Run 3, and Appendix D gives a detailed description of the method discovered by each run., however, reveal that two of the three runs drifted beyond that task: Run 1 toward stability selection and performance weighting, and Run 2 toward a ranking-based method with bias correction. Run 3 stayed closest to the original task, using adaptive LASSO with forward cross validation for lambda selection, though it added an egalitarian elastic net blend. The holdout makes these departures evaluable. Run 2’s bias correction, which exploits serial correlation in combined forecast errors, turns out to generalize and accounts for most of its holdout gain. Run 3’s egalitarian elastic net blend, which drove much of its search-sample improvement, degrades holdout performance. The protocol does not prevent the agent from leaving the intended design space; it ensures that every departure is recorded, so that a holdout or other external evidence can later distinguish structural insight from sample adaptation.
@@ 4 Conclusion
Agent loops can widen researcher degrees of freedom, but they need not do so invisibly. This paper adapts a minimal agent loop architecture to an auditable protocol for agentic specification search and adds a post-search holdout evaluation, requiring only a written instruction contract, a fixed evaluator, a single editable script, a finite search budget, and a complete experiment log. In the forecast combination illustration based on DieboldShin2019, three independent runs improve on the search sample, but a genuine holdout shows that only two of the three discovered methods generalize. The contribution of the protocol is not to eliminate adaptive search, but to make it visible, bounded, and evaluable. Because the protocol disciplines the search process rather than the subject matter, it applies wherever automated specification search is used. Appendix F sketches how the same architecture maps to regression specification search, structural VAR identification, and other standard empirical tasks.
Several limitations deserve emphasis. First, a language model may encode background knowledge that overlaps with the holdout period, so residual look-ahead concerns remain. Second, the model’s internal trade-offs between score improvement and contract boundaries are opaque; it is often unclear how those objectives are balanced. Third, the effective search class may recombine existing ideas more readily than it generates novel ones. Fourth, the current search over $\mathcal{T}(C)$ is greedy: propose one candidate, evaluate it, then keep or discard it. More structured strategies, such as the evolutionary population-based search in AlphaEvolve (NovikovEtAl2025) or tree search in AygunEtAl2025, could explore the space more efficiently. Fifth, the protocol is a practical guardrail against data-adaptive search, but it is not a substitute for formal post-selection inference or other corrections for p-hacking.
@@ A.1 Four-file architecture
program.md is a plain-text instruction contract that defines the research objective, the admissible modifications, the rules, and the experiment loop of Algorithm 1. train.R is the editable script containing one function the agent rewrites. prepare.R is the immutable evaluator: it loads the search sample, sources train.R, and returns the scalar score; the holdout data is entirely outside the agent’s workspace. results.tsv is a tab-separated audit log in which each row records a commit identifier, the score, the outcome status, and a verbal description of the strategy attempted. In the working directory, the git commit history preserves the actual code of every retained $\tau_{k}$. [^3] The replication archive includes the TSV logs and final code but not the full git object store. Full recoverability from commit history requires access to the original working directory.
@@ A.2 Triggering the agent loop
A.3 Isolating the holdout
The post-search holdout $D^{H}$ must remain entirely outside the agent’s workspace during the search. In our application, the holdout period (2017Q1–2025Q4) was constructed after the agentic search was complete, so no holdout data existed in the working directory while the agent was running. More generally, researchers should ensure that holdout data is not accessible to the agent during the search. Because current coding agents can access the internet and execute arbitrary code, an agent could in principle download or construct holdout period data on its own. Disabling internet access during the search, restricting the agent’s file system permissions, or simply withholding the holdout data from the working directory are practical safeguards. Without such boundaries, the transparency of the holdout evaluation cannot be guaranteed.
@@ E.3 Full holdout results
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) and best individual use parameters selected on the search sample and are evaluated without re-optimization on the holdout. peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately.
@@ E.5 Independent dataset robustness
  • Notes: Same as Table 4 but using the independently constructed dataset (all rows from downloaded SPF and Eurostat data, no verbatim copy from the original H1_gdp.csv). Search-sample RMSEs differ from Table 4 due to GDP vintage differences and independent imputation. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test using the EWC fixed-$b$ approximation (ShinSchor2026); see Table 4 notes for details and caveats. peLASSO ex post (original) and best individual use parameters selected on the search sample; see Table 4 notes for details.
@@ Appendix F Illustration: forecast combination and beyond
A post-search holdout can also help reduce concerns about p-hacking, but only in a limited sense. Its value is that it separates discovery from evaluation: the agent may search adaptively over many specifications on the search sample, but the selected output is then judged on data not used during that search. This makes sample-specific improvements less likely to be mistaken for robust findings. At the same time, a holdout is not a complete solution. If the holdout is consulted repeatedly and the method is revised in response, it becomes part of the search process itself. And in inference-heavy settings, holdout performance on a diagnostic criterion does not by itself validate identification, causal interpretation, or post-selection inference. The holdout should therefore be understood as a practical guardrail against hidden adaptive search, not as a full cure.